# Two point functions and quantum fields in the anti-de Sitter universe [^1] Dedicated to the memory of Henri Epstein

### v1

@@ 13.1 The other cycle
In preparation for the general case, it is instructive to integrate also over the other non-compact parabolic cycle. (see Figure 3). Even though we expect a result identical to (13.174) the way to reach it is rather different as the integration cycle is now unbounded and touches infinity. The delicate point is the behavior of the waves at infinity and the way to circumvent such a difficulty will give crucial indications on how to deal with the problem in general spacetime dimension $d$.
The relevant cycle is
@@
This condition has to be interpreted as follows: in formula (13.176) there are two disconnected semi-infinite intervals of integration; the phase of the second factor of the integrand has to be chosen with two different determinations. Eq. (13.176) actually means
@@
$=$ $\frac{\Gamma(-\lambda)\sin(\pi\lambda)\sqrt{u^{\prime}}}{4\pi\sqrt{2\pi}}\int_{-\infty}^{0}e^{i\omega(t-t^{\prime}-is^{\prime})}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}\left(|\omega|u^{\prime}\right)}{}d\omega\int_{u}^{\infty}\cos(\omega\chi-\pi\lambda)\left(\frac{\eta^{2}-u^{2}}{2u}\right)^{\lambda}d\chi$ (13.186)
@@
As anticipated, we recover Eq. (13.174), though by means of a significantly more involved procedure. In the case of general dimensions, the two separate integration cycles contribute differently. Therefore, it is essential to account for both contributions to obtain the complete result.
@@ 14 Waves and two-point functions in Poincaré coordinates: general case
Let us now face the problem of diagonalizing the two-point function (1.1) in Poincaré coordinates in arbitrary spacetime dimension $d$. To this end let us choose $x_{1}=x_{1}({\rm x},u)$ real and $z_{2}=z_{2}({\rm z}^{\prime},u^{\prime})\in{\cal Z}_{+}$ with ${\rm z}^{\prime}\in T_{+}$ in the relevant complex Minkowskian brane.
The integration cycle consists of two disconnected parts, denoted as $\gamma_{P}(x_{1})={\gamma^{+}_{P}(x_{1})\,\cup\,\gamma^{-}_{P}(x_{1})}$ where
$\gamma^{+}_{P}(x_{1})=\left\{\xi(\eta)\in\gamma_{P}^{+}:\left(\eta^{0}-{\rm x}^{0}\right)^{2}\leq\left(\vec{\eta}-\vec{\rm x}\right)^{2}+u^{2}\right\},$ (14.188)
$\gamma^{-}_{P}(x_{1})=\left\{\xi(\chi)\in\gamma_{P}^{-}:\left(\chi^{0}+{\rm x}^{0}\right)^{2}\geq\left(\vec{\chi}+\vec{\rm x}\right)^{2}+u^{2}\right\}.$ (14.189)
As emphasized previously, unlike the two-dimensional case, it is essential to integrate over both disconnected cycles. The following expression is the one that requires evaluation.
@@
where the first line is the integral over $\gamma^{+}_{P}(x_{1})$} and the second two lines are the integral over $\gamma^{-}_{P}(x_{1})$}. The challenging task of evaluating the three integrals above will be rewarded by an exceptionally simple final result. Additionally, the calculation will yield new integral formulae involving Bessel functions.
Let us start with the integral over $\gamma^{+}_{P}(x_{1})$}; in this case the relevant Fourier transform of the waves is as follows:
$\widetilde{\phi^{\pm}_{\lambda}}({\rm k})=\int(x(t\pm i\epsilon,\vec{\rm x},u)\cdot\xi(\eta))^{\lambda}e^{i{\rm k}{\rm x}}\,d{\rm x}\,=e^{i{\rm k}\eta}\int\left[\frac{u}{2}-\frac{({\rm y}^{0}\pm i\epsilon)^{2}-\vec{\rm y}^{2}}{2u}\right]^{\lambda}e^{i{\rm k}{\rm y}}d{\rm y}$ (14.195)
where ${\rm y}={\rm x}-\eta$ (and we left the dependence on $u$ of the lhs implicit). Let us chose ${\rm k}=({\rm k}^{0},\vec{0})$ and perform the radial integration first:
$\widetilde{\phi^{+}_{\lambda}}({\rm k}^{0},0)=e^{i{\rm k}^{0}\eta^{0}}\frac{2\pi^{\frac{d-2}{2}}}{\Gamma(\frac{d-2}{2})}\int_{-\infty}^{\infty}d{\rm y}^{0}\int_{0}^{\infty}\left[\frac{u}{2}-\frac{({\rm y}^{0}+i\epsilon)^{2}-r^{2}}{2u}\right]^{\lambda}e^{i{\rm k}^{0}{\rm y}^{0}}r^{d-3}\,dr=$ (14.196)
@@
valid for $\mathrm{Re}\,(2\lambda+d)<2$ and $\mathrm{Re}\,d>2$ (these restrictions will be removed at the end when the integral will be considered in the sense of distributions).
When ${\rm k}^{0}>0$ the contour of integration may be pushed to infinity and the integral is shown to be zero. The geometric and physical reason behind this possibility is once more the already mentioned immersion of the Minkowskian tubes of the branes in the AdS chiral tuboids.
@@
By comparing Eqs. (14.198) and (14.199) we deduce the identity
$\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{1}{4}+\frac{\lambda}{2}}J_{-\frac{1}{2}-\lambda}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)=2^{\frac{d-4}{2}}\Gamma\left(\frac{d-2}{2}\right)|{\rm k}^{0}|^{1-\frac{d}{2}}u^{\frac{d-1}{2}+\lambda}J_{-\frac{d-1}{2}-\lambda}(|{\rm k}^{0}|u)$ (14.200)
valid when the integral converges. This identity is special instance of a known integral formula by Watson [21] and will be used later.
By restoring Lorentz invariance we obtain the Fourier integral representation of the wave (which we write in a way ready to be used in our calculation):
@@
where the factor $\theta(-{\rm k}^{0})$ guarantees the integral convergence as long as ${\rm z}^{\prime}\in T_{+}$.
We are now ready to evaluate the first line (14.194) (i.e. the integral over the cycle $\gamma^{(+)}_{P}(x_{1})$). Without loss of generality we may simplify the calculation by choosing ${\rm x}=0$ i.e. $x_{1}=x_{1}(0,u).$ Proceeding as in the two-dimensional case (see Eq. (13.172)) we are led to examine the integral
$=\frac{2\sqrt{2}\pi^{\frac{d-1}{2}}|{\rm k}^{0}|^{-\lambda-\frac{1}{2}}\Gamma(\lambda+1)u^{-\lambda}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{\lambda}{2}+\frac{1}{4}}J_{\lambda+\frac{1}{2}}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)dr$ (14.205)
where we set $|\vec{\eta}|=r$ and integrated over the angles; we have assumed $\mathrm{Re}\,\lambda>-1$.
Though similar to (14.199) the integral on the rhs of (14.206) seems to be unknown. Following the example of Eq. (14.199), we may evaluate it by integrating first the variable $r$ on the lhs of (14.206):
@@
By inserting Eqs. (14.203) and (14.214) into Eq. (14.194) we get the first part of the result computed in the special case when $x_{1}=x_{1}(0,u)$:
@@
Let us now focus on the other cycle $\gamma^{(-)}_{P}(x_{1})$. Here the situation is more complicated because the cycle goes through infinity and, as in Eq. (13.179), we need the Fourier transforms of the waves whose phase is vanishing at infinity. The analogous of Eq. (13.177) is here
@@
$=$ $-\frac{4\pi\cos\left(\frac{\pi d}{2}\right)\Gamma(\lambda+1)\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{\Gamma(d+\lambda-1)}\int e^{-i{\rm k}({\rm x}-x^{\prime})}\theta({\rm k}^{0})\theta({\rm k}^{2})J_{\frac{d-1}{2}+\lambda}(u\sqrt{k^{2}})J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{k^{2}}\right)$ (14.235)
@@
The remarkable formula (14.240) shows that the maximally analytic AdS two-point function can be represented as a Källén-Lehmann superposition of maximally analytic two-point functions defined in a Minkowski space with one dimension less, explicitly linking the analytic and geometric structures of AdS and Minkowski spacetimes. Eq. (14.240) was anticipated in [23, 24] by employing Hilbert space methods suitable for generic warped manifolds; the value of the derivation in [23, 24] was also to clarify the Breitenlohner and Friedmann phenomenon [9] in terms of the self-adjointness of a one-dimensional transverse Schrodinger operator; however, a crucial part of the proof was based on a known Hankel transform of the Legendre function $Q$. The present derivation is fully self-contained and specifically tailored to the AdS geometry.
@@ 15 Green functions, Schwinger functions and Wick rotations
Eq. (14.240) may be used to promptly build a new explicit integral of the Schwinger function (8.80) extended to non coincident Euclidean points of $EAdS_{d}$:
@@
There exists a commonly used integral representation of the AdS Schwinger functions called “split representation” [25, 26, 27, 28]; it may be obtained by first representing the Legendre function of the second kind $Q$ as a Mehler-Fock transform of the corresponding Legendre function of the first kind $P$ [19, 29] and then expanding $P$ in plane waves on the Lobachevsky manifold [17, 30]. The advantage of the representation (15.249) is that it has admits a counterpart on the reals that may be obtained by Wick rotation while the split representation cannot be Wick-rotated. Here is the corresponding representation of the Feynman propagator of the Klein-Gordon operator $\Box+m^{2}_{\lambda}$ in the real manifold $AdS_{M}$:

### v2

@@ 13.1 The other cycle
In preparation for the general case, it is instructive to integrate also over the other non-compact parabolic cycle (see Figure 3). Even though we expect a result identical to (LABEL:24), the method to reach it is rather different, as the integration cycle is now unbounded and extends to infinity. The delicate point lies in the behavior of the waves at infinity, and the strategy to circumvent this difficulty will provide crucial guidance for handling the problem in general spacetime dimension $d$.
The relevant relative cycle is
@@
This condition is to be interpreted as follows. In formula (13.188), there are two disconnected semi-infinite intervals of integration; the phase of the second factor of the integrand must be chosen with two different determinations, one for each connected part of the integration cycle. Equation (13.188) thus actually means
@@
$=$ $\frac{\Gamma(-\lambda)\sin(\pi\lambda)\sqrt{u^{\prime}}}{4\pi\sqrt{2\pi}}\int_{-\infty}^{0}e^{i\omega(t-t^{\prime}-is^{\prime})}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}\left(|\omega|u^{\prime}\right)}{}d\omega\times$ (13.199)
${}\ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\int_{u}^{\infty}\cos(\omega\chi-\pi\lambda)\left(\frac{\eta^{2}-u^{2}}{2u}\right)^{\lambda}d\chi=$
@@
As anticipated, we recover the same Hankel representation (LABEL:24), although by a significantly more involved procedure. In general spacetime dimension $d>2$, the two separate integration cycles contribute differently. Therefore, it is essential to account for both contributions in order to obtain the complete result.
@@ 14 Waves and two-point functions in Poincaré coordinates: general case
Let us now face the problem of diagonalizing the two-point function (1.1) in Poincaré coordinates in arbitrary spacetime dimension $d$. To this end, we choose
$x_{1}=x_{1}({\rm x},u)\in AdS_{M}\quad\text{and}\quad z_{2}=z_{2}({\rm z}^{\prime},u^{\prime})\in{\cal Z}_{+},\quad{\rm z}^{\prime}\in T_{+},$
in the relevant complex Minkowskian brane. The integration cycle consists of two disconnected parts, denoted as
$\gamma_{P}(x_{1})=\gamma_{P}^{+}(x_{1})\,\cup\,\gamma_{P}^{-}(x_{1}),$
where
$\gamma_{P}^{+}(x_{1})$ $=$ $\left\{\xi(\eta)\in\gamma_{P}^{+}:(\eta^{0}-{\rm x}^{0})^{2}\leq(\vec{\eta}-\vec{{\rm x}})^{2}+u^{2}\right\},$ (14.201)
$\gamma_{P}^{-}(x_{1})$ $=$ $\left\{\xi(\chi)\in\gamma_{P}^{-}:(\chi^{0}+{\rm x}^{0})^{2}\geq(\vec{\chi}+\vec{{\rm x}})^{2}+u^{2}\right\}.$ (14.202)
As emphasized previously, unlike the two-dimensional case, it is essential to integrate over both disconnected cycles. The following expression is the one that requires evaluation:
@@
Here, the first line corresponds to the integral over $\gamma_{P}^{+}(x_{1})$, while the second and third lines correspond to the integral over $\gamma_{P}^{-}(x_{1})$. The challenging task of evaluating these three integrals will be rewarded by an exceptionally simple final result. Moreover, the calculation will yield new integral formulae involving Bessel functions.
Let us start with the integral over $\gamma_{P}^{+}(x_{1})$. In this case, the relevant Fourier transform of the waves is given by
$\widetilde{\phi^{\pm}_{\lambda}}({\rm k})$ $=$ $\int\big(x({\rm x}^{0}\pm i\epsilon,\vec{{\rm x}},u)\cdot\xi(\eta)\big)^{\lambda}\,e^{i{\rm k}\cdot{\rm x}}\,d{\rm x}$ (14.208)
$=$ $e^{i{\rm k}\cdot\eta}\int\left[\frac{u}{2}-\frac{({\rm y}^{0}\pm i\epsilon)^{2}-\vec{{\rm y}}^{2}}{2u}\right]^{\lambda}e^{i{\rm k}\cdot{\rm y}}\,d{\rm y},$ (14.209)
where ${\rm y}={\rm x}-\eta$, and we leave the dependence on $u$ of the left-hand side implicit.
Let us choose ${\rm k}=({\rm k}^{0},{0})$ and perform the radial integration first:
$\widetilde{\phi^{+}_{\lambda}}({\rm k}^{0},0)$ $=$ $e^{i{\rm k}^{0}\eta^{0}}\frac{2\pi^{\frac{d-2}{2}}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{-\infty}^{\infty}d{\rm y}^{0}\int_{0}^{\infty}\left[\frac{u}{2}-\frac{({\rm y}^{0}+i\epsilon)^{2}-r^{2}}{2u}\right]^{\lambda}e^{i{\rm k}^{0}{\rm y}^{0}}r^{d-3}\,dr$ (14.210)
@@
which is valid for $\mathrm{Re}\,(2\lambda+d)<2$ and $\mathrm{Re}\,d>2$. These restrictions will be removed at the end when the integral will be considered in the sense of distributions.
When ${\rm k}^{0}>0$, the contour of integration can be pushed to infinity, and the integral is then seen to vanish. The geometric and physical reason behind this possibility is, once again, the previously mentioned immersion of the Minkowskian tubes of the branes into the AdS chiral tuboids.
@@
By comparing Eqs. (14.212) and (14.213), we deduce the identity
$\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{1}{4}+\frac{\lambda}{2}}J_{-\frac{1}{2}-\lambda}\Big(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\Big)\,dr=2^{\frac{d-4}{2}}\Gamma\Big(\frac{d-2}{2}\Big)|{\rm k}^{0}|^{\frac{2-d}{2}}u^{\frac{d-1}{2}+\lambda}J_{-\frac{d-1}{2}-\lambda}(|{\rm k}^{0}|u),$ (14.214)
valid when the integral converges. This identity is a special case of a known integral formula by Watson [19] and will be used later.
Restoring Lorentz invariance, we obtain the Fourier integral representation of the wave (written in a form ready to be used in our calculation):
@@
where the factor $\theta(-{\rm k}^{0})$ guarantees convergence of the integral as long as ${\rm z}^{\prime}\in T_{+}$.
We are now ready to evaluate the first line of Eq. (14.207) (i.e., the integral over the cycle $\gamma_{P}^{+}(x_{1})$). Without loss of generality, we may simplify the calculation by choosing ${\rm x}=0$, i.e.,
$x_{1}=x_{1}(0,u).$
Proceeding as in the two-dimensional case (see Eq. (13.184)), we are led to examine the integral
$\frac{2\pi^{\frac{d-2}{2}}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}\,dr\int_{-\sqrt{r^{2}+u^{2}}}^{\sqrt{r^{2}+u^{2}}}e^{i{\rm k}^{0}\eta^{0}}\left(\frac{u^{2}+r^{2}-(\eta^{0})^{2}}{2u}\right)^{\lambda}d\eta^{0}=$
$=\frac{2\sqrt{2}\pi^{\frac{d-1}{2}}|{\rm k}^{0}|^{-\lambda-\frac{1}{2}}\Gamma(\lambda+1)u^{-\lambda}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{\lambda}{2}+\frac{1}{4}}J_{\lambda+\frac{1}{2}}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)dr.$ (14.218)
Here we set $|\vec{\eta}|=r$ and integrate over the angular variables, assuming that $\mathrm{Re}\,\lambda>-1$. Although the integral on the right-hand side of (14.219) is similar to that in (14.213), it does not appear to be directly known. Following the procedure used in Eq. (14.213), we evaluate it by first integrating over the variable $r$ on the left-hand side of (14.219):
@@
By substituting Eqs. (14.217) and (14.227) into Eq. (14.207), we obtain the first part of the result, computed for the special case in which $x_{1}=x_{1}(0,u)$:
@@
Let us now focus on the other cycle $\gamma^{(-)}_{P}(x_{1})$. In this case the situation is more complicated because the cycle passes through infinity and, as in Eq. (13.191), we need the Fourier transforms of the waves whose phase vanishes at infinity. The analogue of Eq. (13.189) is here
@@
$=$ $-\frac{4\pi\cos\left(\frac{\pi d}{2}\right)\Gamma(\lambda+1)\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{\Gamma(d+\lambda-1)}\int e^{-i{\rm k}({\rm x}-{\rm x}^{\prime})}\theta({\rm k}^{0})\theta({\rm k}^{2})J_{\frac{d-1}{2}+\lambda}(u\sqrt{k^{2}})J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{k^{2}}\right)$ (14.248)
@@
The remarkable Hankel formula (14.253) shows that the maximally analytic AdS two-point function can be represented as a Källén-Lehmann superposition of maximally analytic two-point functions defined in a Minkowski space with one dimension less, explicitly linking the analytic and geometric structures of AdS and Minkowski spacetimes.
Eq. (14.253) was anticipated in [27] and rediscovered by [29, 30]. In [25, 26] it was derived by employing Hilbert space methods suitable for generic warped manifolds; the value of the derivation in [25, 26] was also to clarify the Breitenlohner and Friedmann phenomenon [9] in terms of the self-adjointness of a one-dimensional transverse Schrodinger operator; however, a crucial part of the proof was based on a known Hankel transform of the Legendre function $Q$ [17]. The present derivation is fully self-contained and specifically tailored to the geometry of the complex AdS manifold and the AdS spectral condition.
@@ 15 Green functions, Schwinger functions and Wick rotations
Eq. (14.253) may be used to promptly build a new explicit Hankel representation of the Schwinger function (8.91) extended to non coincident Euclidean points of $EAdS_{d}$:
@@
There exists a commonly used integral representation of the AdS Schwinger functions called “split representation” [34, 35, 36, 37]; it may be obtained by first representing the Legendre function of the second kind $Q$ as a Mehler-Fock transform of the corresponding Legendre function of the first kind $P$ [18, 38] and then expanding $P$ in plane waves on the Lobachevsky manifold [14, 39].
The advantage of the representation (15.262) is that it has admits a counterpart on the reals that may be obtained by Wick rotation while the split representation cannot be Wick-rotated (see appendix B). Here is the corresponding representation of the Feynman propagator of the Klein-Gordon operator $\Box+m^{2}_{\lambda}$ in the real manifold $AdS_{M}$: